# Higgs Inflation Model with Small Non-Minimal Coupling Constant

### v4

@@ 3.1 About the shape of the TMT-effective potential and other features of the model
With the obtained value of the parameter $\lambda$, eq.(3.25), we return to the discussion of the shape of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, eqs.(3.6) and (3.7), the graph of which in the entire region $\varphi\geq 0$ is shown [^10] Remember that we chose $q^{4}=3\cdot 10^{-10}$ which ensures the TMT potential with one plateau. in figure 2. Two graphs of $\zeta(\varphi)$ for the case of $X_{\varphi}=0$ are also presented in figure 2; for these two graphs, in addition to the parameters listed above, two corresponding values of $\delta_{p}=0.1$ and $\delta_{p}=10^{-8}$ are chosen in the definition $b_{p}=0.5(1+\delta_{p})$ given by eq.(3.2). As will be explained in section 5, for any admissible values of $\varphi$ the function $\zeta$ must be positive. Therefore, from the shape of curve 2 in figure 2 it follows that the choice $\delta_{p}=0.1$ is unsatisfactory. It should be noted that the graph of $\zeta(\varphi)$ for $\delta_{p}=10^{-8}$, represented by curve 3 in figure 2, is to some extent illustrative, since the case $X_{\varphi}=0$ is unrealistic in general. However, as shown in ref. [22], in the regime of slow-roll inflation, which occurs on the plateau part of $U_{eff}^{(tree)}(\varphi)$, i.e., for $\varphi>6M_{P}$, the contribution of $X_{\varphi}$ to the constraint (3.1) is bounded from above by the following estimate $\frac{X_{\varphi}}{M_{P}^{4}}\approx 128\lambda\cdot e^{-2\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}<9\cdot 10^{-18}$. Therefore, for $\varphi>6M_{P}$, the curve 3 in figure 2 can be considered as quite reliable graph for $\zeta$. In the opposite region, where $\varphi\lll M_{P}$, one can expect that $X_{\varphi}\lll M_{P}^{4}$ and therefore the $\zeta(\varphi)$ graph is also reliable. However, it is quite natural to suppose that soon after the end of inflation the contribution of $X_{\varphi}$ to the constraint (3.1) may become significant. Nevertheless, from the graph with $\delta_{p}=10^{-8}$ it follows that the decrease in $\varphi$ at the inflationary stage is accompanied by a monotonous increase in $\zeta(\varphi)$ from positive values close to zero [^11] The fundamental significance of the condition $\zeta>0$ will be the subject of study in section 5.; and for $\varphi\lll M_{P}$, with a decrease in $\varphi$, $\zeta(\varphi)$ also increases monotonically, asymptotically approaching the value $\zeta=\zeta_{v}=1$.
@@
The representation of $U_{eff}^{(tree)}(\varphi)$ in figure 2 does not allow to see important details of the shape, which differ greatly in the regions $\varphi>M_{P}$ and $\varphi\ll M_{P}$. This is due to the greatly different behaviour of the factor $F(\varphi)$, eq.(3.7), in the regions $\varphi>M_{P}$ and $\varphi\ll M_{P}$. The analysis simplified if we use the value of the primordial model parameter $m\approx 0.7GeV$ obtained in ref. [22], that is, $\frac{m^{2}}{M_{P}^{2}}\approx 8\cdot 10^{-38}$. Then the form of the function $F(\varphi)$ depends on the ratio of the second and fifth terms in the numerator of $F(\varphi)$. A direct estimate shows that the value $\varphi\approx 1.7\cdot 10^{-4}M_{P}\approx 4.1\cdot 10^{14}GeV$ divides the entire region $\varphi\geq 0$ into two intervals, where one of the indicated terms dominates over the other. Accordingly, we arrive at the following asymptotic representation for the TMT-effective potential in two opposite regions
@@
$U_{eff}^{(tree)}(\varphi)=\begin{cases}\frac{\lambda M_{P}^{4}}{8\xi^{2}}{\tanh}^{4}z\cdot\frac{3q^{4}+\frac{\lambda}{8\xi^{2}}\sinh^{4}z}{4q^{4}+\frac{\lambda}{8\xi^{2}}\sinh^{4}z}&\text{for}\quad\varphi\gg 10^{-4}M_{P}\\ \frac{\zeta_{v}+b_{p}}{(1+\zeta_{v})^{2}}\frac{\lambda}{4}\phi^{4}-\frac{\zeta_{v}-b_{p}}{(1+\zeta_{v})^{2}}\frac{m^{2}}{2}\phi^{2}&\text{for}\quad\phi\ll 10^{-4}M_{P},\end{cases}$ (4.1)
@@
The result in region $\phi\ll 10^{-4}M_{P}$, presented in figure 3, will be more understandable if we consider the expression for the TMT-effective Lagrangian $L_{eff}^{(tree)}(\varphi,X_{\varphi})$ given by eqs.(3.4)-(3.10). Considering that, as follows from eqs.(3.9) and (3.10)
@@
In the region $\phi\ll 10^{-4}M_{P}$ it is natural to assume that $|X_{\phi}|\lesssim|U_{eff}(\phi)|_{\phi\ll 10^{-4}M_{P}}|\ll q^{4}M_{P}^{4}\sim(10^{16}GeV)^{4}$. Then it is easy to see that the second term in (4.3) is many orders of magnitude smaller than the first. In this approximation, the field $\phi$ equation of motion, obtained using the Lagrangian $L_{eff}^{(tree)}|_{\phi\ll 10^{-4}M_{P}}$ and then dividing by a constant factor $\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}$, coincides with the $\phi$-equation near the vacuum, described by eqs.(2.29) and (2.30), found in section 2.3. This analysis allows us to get some idea of what structure the field equations might have for $\phi\lesssim 10^{-4}M_{P}$.
@@ 5 Constraint, orientability of the space-time manifold
and their impact on cosmology
@@ 5.2 Constraint, initial conditions
Let’s look at the implications that can be drawn from the inequalities (5.3) with the model parameters used in figure 2, where we chosed $q^{4}=3\cdot 10^{-10}$. The shape of the potential $U_{eff}^{(tree)}(\varphi)$ shows that, with the exception of the very end, inflation can occur at $\varphi\gtrapprox 6M_{P}$; if evolution starts from other scenario, possibility of which will be found in sec.6, it is natural to assume that this also occurs at $\varphi\gtrapprox 6M_{P}$. The height of the potential plateau is between $10^{-10}M_{P}^{4}\lesssim U_{eff}(\varphi)\lesssim 1.04\cdot 10^{-10}M_{P}^{4}$. The latter will be used in the numerical estimates made below. The remaining free parameter is $10<\delta_{p}<$ 1 and it turns out that changing its value can lead to qualitatively different results. As numerical examples, consider two cases $\delta_{p}=0.1$ and $\delta_{p}=10^{-8}$, corresponding to two graphs $\zeta(\varphi)$ for $X_{\varphi}=0$ in figure 2.
@@
    $-5.3\cdot 10^{-13}\lesssim\frac{X_{\varphi}}{M_{P}^{4}}\lesssim\frac{\lambda}{8\xi^{2}}\approx\frac{U_{eff}^{(tree)}(\varphi)}{M_{P}^{4}}\approx 10^{-10}$ (5.7)
@@
    $-5.3\cdot 10^{-13}M_{P}^{4}\lesssim\rho_{kin}^{(in)}-\rho_{grad}^{(in)}\lesssim U_{eff}^{(tree)}(\varphi_{(in)})\approx 10^{-10}M_{P}^{4}$ (5.8)
@@
Thus, we come to the conclusion that there is a quite wide range of model parameters in which conditions (1.1) are satisfied. At this stage, it is worth noting once again that the described results follow only from the condition $\zeta>0$, which, as discussed above, means that the space-time manifold of our Universe is orientable. Recall also that, as explained in the paragraph after eq.(2.17), the origin of the restrictions imposed by the constraint on the dynamics of the field $\phi$ are in no way dependent on the dynamics of the field $\phi$; this is true in general and also in this particular case with respect to the condition $\zeta>0$. However, as noted at the beginning of this section, in order to assert that the conditions necessary for the onset of inflation are met, we must take into account the influence of the K-essence type structure.
@@ 6.1 General overview and possible scenarios
Using the TMT-effective action, eqs.(3.4)-(3.10), in the spatially flat FLRW universe, the following Friedmann and scalar field equations are obtained after using eqs.(3.11), (3.4), (3.5):
$H^{2}=\frac{1}{3M_{P}^{2}}\Bigl{[}\Big{(}K_{1}-\frac{3}{2}\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Big{)}\cdot\frac{1}{2}\dot{\varphi}^{2}+U_{eff}^{(tree)}(\varphi)\Bigr{]},$ (6.1)
$\Bigl{(}K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr{)}\ddot{\varphi}+\Bigl{(}K_{1}-\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr{)}3H\dot{\varphi}+\Bigl{(}K_{1}^{\prime}-\tilde{K}^{\prime}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr{)}\dot{\varphi}^{2}+U_{eff}^{(tree)^{\prime}}(\varphi)=0.$ (6.2)
@@
It is obvious that the structure of the cosmological equations (6.1) and (6.2) differs significantly from the usual ones. Indeed, first of all, in the place where the kinetic energy density should be in the energy density, we find an expression
@@
from where we get the upper bound
@@
$\ddot{\varphi}+f_{1}(y)3H\dot{\varphi}+f_{2}(y)\frac{\dot{\varphi}^{2}}{2}+\frac{U_{eff}^{(tree)\prime}(\varphi)}{K_{1}-3\tilde{K}_{2}\cdot y}=0,$ (6.10)
@@
$y=\frac{\frac{1}{2}\dot{\varphi}^{2}\cdot 10^{-10}}{U_{eff}^{(tree)}}$ $K_{eff}$ $f_{1}(y)$ $\frac{U_{eff}^{(tree)^{\prime}}}{K_{1}-3\tilde{K}_{2}y}$
@@
$0.80.52<y<$ 0.8 $K_{eff}<0$ $f_{1}(y)>0$wrong" $w<-1$
$0.520.35<y<$ 0.52 $K_{eff}<0$ $f_{1}(y)<0$wrong" $w>-1$
@@
The graph of EOS is presented in figure 5, from which it is evident that the evolution of the Universe can begin with $w<-1$. Thus, using the terminology adopted in the models of present-day evolution of the Universe, we come to the conclusion that the classical evolution of the Universe can begin with the stage of phantom dynamics.
@@
Figure 5: Plot of the function (6.16) describing the $\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}$-dependence of the equation of state $w=\frac{p}{\rho}$ in the region $\varphi>6M_{P}$, where the TMT effective potential has a plateau $U_{eff}(\varphi)\approx 10^{-10}M_{P}^{4}$. The condition (6.8), necessary for the total initial energy density to be positive, is manifested in the fact that the line $y=\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\cdot 10^{10}=0.8$ is a vertical asymptote for the function $w(y)$ and $w\to-\infty$ for $y\to 0.8^{-}$. The results presented in eqs.(6.12)-(6.14), together with the shape of the graph $w(y)$, require dividing the interval $0.80<y<$ 0.8 into four regions with different types of dynamics: 1) region $0.80.52<y<$ 0.8 - pathological type I with EoS $w<-1$, i.e. phantom dynamics; 2) region $0.520.35<y<$ 0.52 - pathological type I dynamics with EoS $w>-1$; 3) region $0.350.17<y<$ 0.35 - pathological type II dynamics; 4) region $0.170<y<$ 0.17 - normal dynamics.
@@
At the end of this section, at least brief comments on pathological dynamics are appropriate. It would be interesting to study possible solutions with such initial conditions. Preliminary analysis shows that the pathological stages of evolution are dynamically isolated from the stage with normal dynamics. It may also be noted that such an initial pathological stage of the evolution of the Universe could be responsible for the unknown physics, whose existence has been suggested in the interpretation of the BGV theorem[38]. However, in this paper we will limit ourselves to posing these questions and will study them elsewhere.

### v5

@@ 3.1 About the shape of the TMT-effective potential and other features of the model
With the obtained value of the parameter $\lambda$, eq.(3.25), we return to the discussion of the shape of the TMT-effective potential $U_{eff}^{(tree)}(\varphi)$, eqs.(3.7) and (3.8), the graph of which in the entire region $\varphi\geq 0$ is shown [^10] Remember that we chose $q^{4}=3\cdot 10^{-10}$ which ensures the TMT potential with one plateau. in figure 2. Two graphs of $\zeta(\varphi)$ for the case of $X_{\varphi}=0$ are also presented in figure 2; for these two graphs, in addition to the parameters listed above, two corresponding values of $\delta_{p}=0.1$ and $\delta_{p}=10^{-8}$ are chosen in the definition $b_{p}=0.5(1+\delta_{p})$ given by eq.(3.2). As will be explained in section 5, for any admissible values of $\varphi$ the function $\zeta$ must be positive. Therefore, from the shape of curve 2 in figure 2 it follows that the choice $\delta_{p}=0.1$ is unsatisfactory. It should be noted that the graph of $\zeta(\varphi)$ for $\delta_{p}=10^{-8}$, represented by curve 3 in figure 2, is to some extent illustrative, since the case $X_{\varphi}=0$ is unrealistic in general. However, as shown in appendix B, in the regime of slow-roll inflation, which occurs on the plateau part of $U_{eff}^{(tree)}(\varphi)$, i.e., for $\varphi>6M_{P}$, the contribution of $X_{\varphi}$ to the constraint (3.1) is bounded from above by the following estimate $\frac{X_{\varphi}}{M_{P}^{4}}\approx 128\lambda\cdot e^{-2\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}<9\cdot 10^{-18}$. Therefore, for $\varphi>6M_{P}$, the curve 3 in figure 2 can be considered as quite reliable graph for $\zeta$. In the opposite region, where $\varphi\lll M_{P}$, one can expect that $X_{\varphi}\lll M_{P}^{4}$ and therefore the $\zeta(\varphi)$ graph is also reliable. After inflation ends, the contribution of $X_{\varphi}$ to the constraint (3.1) may increase, but we will assume that it does not lead to qualitative changes [^11] The contribution of $X_{\phi}/M_{P}^{4}$ to the constraint (2.35) is negligible if $\frac{\dot{\phi}^{2}}{M_{P}^{4}}\ll 3.6\cdot 10^{-9}$. From the graph with $\delta_{p}=10^{-8}$ it follows that
  • inflation can start from sufficiently large initial values of the scalar field $\varphi\approx 14M_{P}$;
  • the decrease in $\varphi$ at the inflationary stage is accompanied by a monotonous increase in $\zeta(\varphi)$ from positive values very close to zero [^12] The fundamental significance of the condition $\zeta>0$ will be the subject of study in section 5..
For $\varphi\ll M_{P}$, with a decrease in $\varphi$, $\zeta(\varphi)$ also increases monotonically, asymptotically approaching the value $\zeta=\zeta_{v}=1$.
@@
The representation of $U_{eff}^{(tree)}(\varphi)$ in figure 2 does not allow to see important details of the shape, which differ greatly in the regions $\varphi>M_{P}$ and $\varphi\ll M_{P}$. This is due to the greatly different behaviour of the factor $F(\varphi)$, eq.(3.8), in the regions $\varphi>M_{P}$ and $\varphi\ll M_{P}$. The analysis simplified if we use the value of the primordial model parameter $m\approx 0.7GeV$ obtained in ref. [22], that is, $\frac{m^{2}}{M_{P}^{2}}\approx 8\cdot 10^{-38}$. A direct evaluation shows that the entire region $\varphi\geq 0$ can be divided into two opposite regions with the following asymptotic representations for the TMT-effective potential
@@
$U_{eff}^{(tree)}(\varphi)=\begin{cases}\frac{\lambda M_{P}^{4}}{8\xi^{2}}{\tanh}^{4}z\cdot\frac{3q^{4}+\frac{\lambda}{8\xi^{2}}\sinh^{4}z}{4q^{4}+\frac{\lambda}{8\xi^{2}}\sinh^{4}z}&\text{for}\quad\varphi>M_{P}\\ \frac{\zeta_{v}+b_{p}}{(1+\zeta_{v})^{2}}\frac{\lambda}{4}\phi^{4}-\frac{\zeta_{v}-b_{p}}{(1+\zeta_{v})^{2}}\frac{m^{2}}{2}\phi^{2}&\text{for}\quad\phi\ll 10^{-4}M_{P},\end{cases}$ (4.1)
@@
The result in region $\phi\ll 10^{-4}M_{P}$, presented in figure 3, will be more understandable if we consider the expression for the TMT-effective Lagrangian $L_{eff}^{(tree)}(\varphi,X_{\varphi})$ given by eqs.(3.6)-(3.11). An important point is that the kinetic term in $L_{eff}^{(tree)}(\varphi,X_{\varphi})$ is non-canonical. Considering that, as follows from eqs.(3.10) and (3.11)
@@
In the region $\phi\ll 10^{-4}M_{P}$ it is natural to assume that $|X_{\phi}|\lesssim|U_{eff}(\phi)|_{\phi\ll 10^{-4}M_{P}}|\ll q^{4}M_{P}^{4}\sim(10^{16}GeV)^{4}$. Then it is easy to see that the second term in (4.3) is many orders of magnitude smaller than the first. In this approximation, the field $\phi$ equation of motion, obtained using the Lagrangian $L_{eff}^{(tree)}|_{\phi\ll 10^{-4}M_{P}}$ and then dividing by a constant factor $\frac{b_{k}-\zeta_{v}}{1+\zeta_{v}}$, coincides with the $\phi$-equation near the vacuum, described by eqs.(2.31) and (2.32), found in section 2.3.
@@ 5 Constraint, orientability of the space-time manifold
and their impact on the initial coditions
@@ 5.2 Constraint, initial conditions
Let’s look at the implications that can be drawn from the inequalities (5.3) with the model parameters used in figure 2, where we chosed $q^{4}=3\cdot 10^{-10}$. The shape of the potential $U_{eff}^{(tree)}(\varphi)$ shows that, with the exception of the very end, inflation can occur at $\varphi\gtrapprox 6M_{P}$; if evolution starts from other scenario, possibility of which will be found in sec.6, it is natural to assume that this also occurs at $\varphi\gtrapprox 6M_{P}$. The height of the potential plateau is between $10^{-10}M_{P}^{4}\lesssim U_{eff}(\varphi)\lesssim 1.04\cdot 10^{-10}M_{P}^{4}$. The latter will be used in the numerical estimates made below. The remaining free parameter is $0<\delta_{p}<1$ and it turns out that changing its value can lead to qualitatively different results. As numerical examples, consider two cases $\delta_{p}=0.1$ and $\delta_{p}=10^{-8}$, corresponding to two graphs $\zeta(\varphi)$ for $X_{\varphi}=0$ in figure 2.
@@
    $5.3\cdot 10^{-13}\lesssim\frac{X_{\varphi}}{M_{P}^{4}}\lesssim\frac{\lambda}{8\xi^{2}}\approx\frac{U_{eff}^{(tree)}(\varphi)}{M_{P}^{4}}\approx 10^{-10}$ (5.7)
@@
    $5.3\cdot 10^{-13}M_{P}^{4}\lesssim\rho_{kin}^{(in)}-\rho_{grad}^{(in)}\lesssim U_{eff}^{(tree)}(\varphi_{(in)})\approx 10^{-10}M_{P}^{4}$ (5.8)
@@
Thus, we come to the conclusion that there is a quite wide range of model parameters in which conditions (1.1) for the onset of inflation are met . At this stage, it is worth noting once again that the described results follow only from the condition $\zeta>0$, which, as discussed above, means that the space-time manifold of our Universe is orientable. Recall also that, as explained in the paragraph after eq.(2.19), the origin of the restrictions imposed by the constraint on the dynamics of the field $\phi$ are in no way dependent on the dynamics of the field $\phi$; this is true in general and also in this particular case with respect to the condition $\zeta>0$. However, as noted at the beginning of this section, in order to assert that the conditions necessary for the onset of inflation are met, we must take into account the influence of the K-essence type structure.
@@ 6.1 General overview and possible scenarios
Using the TMT-effective potential (3.4) and the TMT-effective action, eqs.(3.5)-(3.11), in the spatially flat FLRW universe, the following Friedmann and scalar field equations are obtained:
$\frac{\dot{a}^{2}}{{a}^{2}}=\frac{1}{3M_{P}^{2}}\Bigl[\Big(K_{1}-\frac{3}{2}\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Big)\cdot\frac{1}{2}\dot{\varphi}^{2}+U_{eff}^{(tree)}(\varphi)\Bigr],$ (6.1)
$\Bigl(K_{1}-3\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr)\ddot{\varphi}+\Bigl(K_{1}-\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr)3\frac{\dot{a}}{a}\dot{\varphi}+\Bigl(K_{1}^{\prime}-\tilde{K}^{\prime}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Bigr)\dot{\varphi}^{2}+U_{eff}^{(tree)^{\prime}}(\varphi)=0.$ (6.2)
@@
The structure of the cosmological equations (6.1) and (6.2) differs significantly from the usual ones. First of all, in the place where the kinetic energy density should be in the energy density, we find an expression
@@
from where we get the upper bound on $\dot{\varphi}_{in}$
@@
$\ddot{\varphi}+f_{1}(y)3\frac{\dot{a}}{a}\dot{\varphi}+f_{2}(y)\frac{\dot{\varphi}^{2}}{2}+\frac{U_{eff}^{(tree)\prime}(\varphi)}{K_{1}-3\tilde{K}_{2}\cdot y}=0,$ (6.10)
@@
$\tilde{y}=10^{10}y=10^{10}\frac{\dot{\phi}^{2}}{2M_{P}^{4}}$ $K_{eff}$ $f_{1}(y)$ $\frac{U_{eff}^{(tree)^{\prime}}}{K_{1}-3\tilde{K}_{2}y}$
@@
$0.52<\tilde{y}<0.8$ $K_{eff}<0$ $f_{1}(y)>0$ "wrong" w<1w<-1
$0.35<\tilde{y}<0.52$ $K_{eff}<0$ $f_{1}(y)<0$ "wrong" w>1w>-1
@@
The graph of EOS is presented in figure 5, from which it is evident that the evolution of the Universe can begin with w<1w<-1. Thus, using the terminology adopted in the models of present-day evolution of the Universe, we come to the conclusion that the classical evolution of the Universe can begin with the stage of phantom dynamics.
@@
Figure 5: Plot of the function (6.16) describing the $\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}$-dependence of the equation of state $w=\frac{p}{\rho}$ in the region $\varphi>6M_{P}$, where the TMT effective potential has a plateau $U_{eff}(\varphi)\approx 10^{-10}M_{P}^{4}$. The condition (6.8), necessary for the total initial energy density to be positive, is manifested in the fact that the line $\tilde{y}=\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\cdot 10^{10}=0.8$ is a vertical asymptote for the function $w(y)$ and $w\to$-$\infty$ for $\tilde{y}\to 0.8^{-}$. The results presented in eqs.(6.12)-(6.14), together with the shape of the graph $w(y)$, require dividing the interval $0<\tilde{y}<0.8$ into four regions with different types of dynamics: 1) region $0.52<\tilde{y}<0.8$ - pathological type I with EOS w<1w<-1, i.e. phantom dynamics; 2) region $0.35<\tilde{y}<0.52$ - pathological type I dynamics with EOS w>1w>-1; 3) region $0.17<\tilde{y}<0.35$ - pathological type II dynamics; 4) region $0<\tilde{y}<0.17$ - normal dynamics.
Figure 6: Three disconnected sets of phase trajectories satisfying eq.(6.10). The set of trajectories between lines A and B corresponds to solutions of the pathological type I dynamics with EOS w<1w<-1. The set of trajectories between lines B and D corresponds to solutions with EOS w>1w>-1, and line C separates the dynamics with the pathological types I and $I$. The set of phase trajectories above line D corresponds to solutions with normal dynsmics. For illustration, for six chosen initial values of ${\phi}_{in}$, three initial values of $\dot{\phi}_{in}$ are chosen respectively such that 1) close to line A (above A); 2) close to line D (below D); 3) close to line D (above D). Additional analysis of the dynamic properties is given in the main text.
Figure 7: Three graphs represent the dependences of $\ln a$ on $\tau=10^{-5}M_{P}t$ for solutions of eqs.(6.1), (6.2) with three initial values of $\dot{\phi}_{in}$ chosen such that 1) close to line A in figure 6 (above A, $z_{in}\equiv\frac{\dot{\phi}_{in}}{M_{P}^{2}}\cdot 10^{5}=-1.24$); 2) close to line D (below D, $z_{in}=-0.61$); 3) close to line D (above D, $z_{in}=-0.57$), while the initial value of ${\phi}_{in}$ is the same for three solutions.
After the presented classification of the types of dynamics depending on the initial values of $\dot{\phi}_{in}$, it is necessary to clarify the characteristic dynamic properties of each of them in the course of cosmological evolution. This can be conveniently done using the phase plane in figure 6, which shows the phase curves corresponding to the initial conditions close to the limiting ones for each type of dynamics. Accordingly, the phase plane is divided by horizontal lines into dynamically disconnected regions. Line A separates the region with $\rho>0$ from the region with $\rho<0$. In the regions below line D, the sign of the active force in eq.(6.10) is "wrong", which is the main reason for the existence of pathological dynamics. Line B separates the regions with opposite sign of the friction term $f_{1}(y)3H\dot{\varphi}$ in eq.(6.10). Phantom-like dynamics with EOS w<1w<-1 can exist only in the region between lines A and B. When the phase curves intersect the dotted line C, the TMT-effective kinetic energy $K_{eff}$ changes sign from $K_{eff}>0$ above C to $K_{eff}<0$ below C. In the region above line D, the active and frictional forces have regular signs, and $K_{eff}>0$. Consequently, the phase curves with initial conditions located above the line D correspond to normal dynamics and, as can be seen, the slow decrease of $\phi$ is accompanied by a rapid approach of $\dot{\phi}$ to zero, which is a characteristic property of the regime of the slow-roll inflation. It is interesting to note that line B is an attractor for the phase curves on both sides, i.e. not only for the phase curves in the region between A and B, but also for the phase curves in the region between D and B.
Three cirves (dotted, dashed and solid) in figure 7 show the cosmic time dependences of the scale factor $a(t)$ for three regions of the phase diagram with the same ${\phi}_{in}$ and three different values of $\dot{\phi}_{in}$: the first (dotted) with $\dot{\phi}_{in}$ slightly above line A , the second (dashed) with $\dot{\phi}_{in}$ slightly below line D , and the third (solid) with $\dot{\phi}_{in}$ slightly above line D. As can be seen, the graphs demonstrate predominantly exponential expansion. The only exception is the initial section of the first, dotted line, which means that for initial conditions close to line A, the initial stage of the cosmological expansion is non-exponential.
@@
At the end of this section, at least brief comments are appropriate. The above analysis shows that pathological regimes of evolution are dynamically isolated from the regime with normal dynamics, which has been found to inevitably describe slow-roll inflation (see also sections 6.2 and 6.3). It may also be noted that the initial pathological stage of the evolution of the Universe predicted by TMSM could be a model describing the unknown physics whose existence was proposed in the interpretation of the BGV theorem[38].